OpenStudy (anonymous):

Solve the inequality. Write your answer in interval notation. Graph. x^2 - 9 / x^2 - 12x + 11 is greater than or equal to zero.

OpenStudy (anonymous):

[(x-3)(x+3)]/[(x-11)(x-1)] >=0 \[(-\infty ,-3]\cup(1,3]\cup(11,+\infty) \]

OpenStudy (anonymous):

wow, thank you.

OpenStudy (anonymous):

interval notation is hard for me

OpenStudy (anonymous):

so on the graph, is the point (0,-3) supposed to point upward?

OpenStudy (anonymous):

it is a solid dot

OpenStudy (anonymous):

is it (3,0) solid dot. (-3,0) solid dot, (1,0) open dot, and (11,0) open dot?

OpenStudy (anonymous):

yes ( ) means open dot and [ ] means closed dot however infinity ( or ) means arrow

OpenStudy (anonymous):

ok. thank you very much